# Hydrogen-Alpha as a Tracer of Star Formation in the SPHINX Cosmological Simulations

### v1

@@
Hydrogen-Alpha as a Tracer of Star Formation in the SPHINX Cosmological Simulations
The Hydrogen-alpha () emission line in galaxies is a powerful tracer of their recent star formation activity. With the advent of JWST, we are now able to routinely observe in galaxies at high redshifts () and thus measure their star-formation rates (SFRs). However, using “classical” SFR() calibrations to derive the SFRs leads to biased results because high-redshift galaxies are commonly characterized by low metallicities and bursty star-formation histories, affecting the conversion factor between the luminosity () and the SFR. In this work, we develop a set of new SFR() calibrations that allow us to predict the SFRs of -emitters at with minimal error. We use the SPHINX cosmological simulations to select a sample of star-forming galaxies representative of the -emitter population observed with JWST. We then derive linear corrections to the classical SFR() calibrations, taking into account variations in the physical properties (e.g., stellar metallicities) among individual galaxies. We obtain two new SFR() calibrations that, compared to the classical calibrations, reduce the root mean squared error (RMSE) in the predicted SFRs by dex and dex, respectively. Using the recent JWST NIRCam/grism observations of -emitters at , we show that the new calibrations affect the high-redshift galaxy population statistics: (i) the estimated cosmic star-formation density decreases by %, and (ii) the observed slope of the star-formation main sequence increases by .
@@
One of the most fundamental and basic properties of galaxies is the rate at which new stars form, the star formation rate (SFR). The SFR is sensitive to the interplay between the gas inflow and outflow rates of galaxies, and the efficiency at which this gas is converted into stars (e.g., Bouché et al., 2010; Lilly et al., 2013; Behroozi et al., 2019). The SFR can fluctuate on different timescales (e.g., Sparre et al., 2015; Iyer et al., 2020; Flores Velázquez et al., 2021), which range from dynamical timescales of molecular clouds (s of Myrs), through galaxy merger timescales ( Myr), to long (Gyr-scale) coherences in accretion rates (e.g., Matthee & Schaye, 2019; Wang & Lilly, 2020; Tacchella et al., 2020; Wan et al., 2025). Observationally, SFRs can be measured using various diagnostics (e.g., Kennicutt & Evans, 2012; De Looze et al., 2014; Davies et al., 2019) that are either based on the light from massive, short-lived stars (e.g., the rest-frame UV), or processed radiation (e.g., nebular lines in HII regions or dust emission). Each of these diagnostics probes different timescales of star formation and is affected by dust attenuation differently.
@@
At high redshifts (), the rest-frame UV emission used to be the most sensitive SFR indicator that could be measured for large galaxy samples, tracing star formation on the Myr timescales (e.g., Bouwens et al., 2015). However, dust corrections in the rest-frame UV are significant, with potentially up to 50% of the star formation rate density at being obscured by dust (e.g., Zavala et al., 2021; Sun et al., 2025). In contrast, the Hydrogen-alpha () emission – an alternative SFR indicator – is emitted in the rest-frame optical ( Å) and thus has a much lower sensitivity to dust attenuation. The luminosity of the line is directly proportional to the luminosity of ionizing photons that are predominantly emitted by very young stellar populations, making sensitive to star formation on shorter timescales ( Myr; e.g., Leitherer et al., 1999). As the JWST can now measure the distribution of luminosities out to based on spectroscopy and narrow-band imaging (e.g., Clarke et al., 2024; Pirie et al., 2024; Covelo-Paz et al., 2025; Fu et al., 2025), it is timely to reassess the role of the H emission as a tracer of the SFR.
@@
The conversion factor between the H luminosity111Throughout this work, we assume a negligible contribution of active galactic nuclei to the H emission. and the SFR primarily depends on assumptions on the properties of stellar populations, in particular the star formation history (SFH), the initial mass function (IMF), binarity and stellar metallicity. The SFH and the IMF are relevant because the ionizing photon emissivity from a single stellar population drops sharply after a few Myr when massive, short-lived O stars leave the main sequence. Binary interactions can extend these ionizing lifetimes (Götberg et al., 2017), smoothing the otherwise steep drop-off after a few Myrs. At fixed stellar mass and age, the metallicity impacts the effective temperature of stellar atmospheres due to line blanketing effects, affecting the ionizing luminosity as well.
@@
The classical SFR() calibration proposed by Kennicutt (1998) (see also Kennicutt & Evans 2012) assumes a constant SFH over Myr and solar metallicity, based on the typical conditions in galaxies in the low-redshift Universe. However, various spectroscopic observations of galaxies at high redshifts suggest that their SFHs are bursty (e.g., Asada et al., 2024; Endsley et al., 2024; Cole et al., 2025; Carvajal-Bohorquez et al., 2025). Given that the gas-phase oxygen abundances are typically measured to be % solar (Curti et al., 2024; Scholte et al., 2025), and early galaxies are expected to be -enhanced due to the rapid formation timescales (e.g. Chruślińska et al., 2024), stellar metallicities are likely to be very low (), implying hotter stars and a more efficient photon production at a fixed SFR (e.g., Theios et al., 2019), compared to low redshift.
@@
The aim of this paper is to revisit the classical SFR() calibrations to predict the SFRs of galaxies at high redshifts (). Rather than using idealistic models with constant SFHs and fixed stellar metallicities, we use simulated galaxies and their modeled luminosities from the SPHINX cosmological radiation-hydrodynamical simulation (Rosdahl et al., 2018; Katz et al., 2023). These cosmological simulations self-consistently model SFHs and metal enrichment processes, and are calibrated to match various observables at high redshifts (). The simulated volume is well matched to the typical volumes of deep JWST surveys, which enables a high resolution that can capture SFR variations on short timescales ( Myr) and approximate the distribution of gas in the interstellar medium relevant for radiative transfer effects.
@@
In Sect. 2, we briefly describe the SPHINX simulations and explain how the H luminosities are derived. In Sect. 3, we present a set of simulation-inferred SFR() calibrations that predict the SFRs of SPHINX galaxies with minimal error. Finally, we discuss the implications of our work in Sect. 4, and give a summary in Sect. 5.
@@ 2 Methods
In this work, we use the simulated galaxy data from SPHINX, a suite of cosmological radiation-hydrodynamic simulations (Rosdahl et al., 2018). SPHINX leverages the adaptive mesh refinement code RAMSES-RT (Rosdahl et al., 2013; Teyssier, 2002) to simulate tens of thousands of galaxies in the epoch of reionization. The ISM is resolved down to co-moving pc, corresponding to the physical resolution of pc at . The dark matter halos are resolved down to the atomic cooling threshold ( M⊙). Hydrodynamics, non-equilibrium thermochemistry, and radiative transfer of Lyman continuum radiation (LyC; Å) are performed self-consistently on-the-fly, with subgrid implementations for gas cooling, star formation, metal enrichment, and feedback. The on-the-fly radiative emission and propagation of LyC photons make SPHINX an excellent model to accurately predict how emission traces star formation in the simulated galaxies.
@@
Our analysis is based on the data from the SPHINX20 Public Data Release, Version 1 (SPDRv1; Katz et al., 2023). SPHINX20 is the largest simulation of the SPHINX suite, with a periodic box with volume of ( cMpc)3 simulated down to a redshift of (Rosdahl et al., 2022). SPHINX20 uses the BPASSv stellar population synthesis (SPS) model (Stanway & Eldridge, 2018) that adopts a Kroupa (2001a) IMF with an upper mass cutoff of 100 M⊙ and a slope of from to M⊙ and from to M⊙. The intrinsic Balmer line luminosities are generated by interpolating the CLOUDY v17.03 models (Ferland et al., 2017) for gas cells that host stellar particles with unresolved Stromgren spheres, or by using the non-equilibrium ionizing fractions from the simulation otherwise. The equivalent widths are calculated by taking into account both the stellar and nebular continuum emission, generated as detailed in Katz et al. (2023). Post-processing of line and continuum radiation, subject to absorption and scattering by dust, is carried out using the radiative transfer code RASCAS (Michel-Dansac et al., 2020) assuming the Small Magellanic Cloud (SMC) dust model from Michel-Dansac et al. (2020) (for details, see Katz et al., 2023).
@@
SPDRv1 only includes SPHINX20 galaxies with M⊙ yr-1, thus excluding objects that are likely too faint to be observed with JWST (Katz et al., 2023). However, an SFR cut leads to incompleteness in the SFR— relation at low luminosities which can bias our results. Therefore, we limit our selection of galaxies from SPDRv1 to those that have an observed luminosity () greater than erg s-1. This limit is motivated by the recent LF measurements from JWST NIRCam/grism that reach similar luminosities at the faint end (e.g., Fu et al., 2025). Crucially, this selection yields a complete -emitter sample, which we check by running RASCAS on all SPHINX20 haloes whose intrinsic luminosity () exceeds the luminosity cut. We also try erg s-1 and erg s-1 cuts, finding little impact on our results. With this additional luminosity cut, our final sample comprises , 66, 48, 36, 22, 17, and 13 galaxies at , 5, 6, 7, 8, 9, and 10, respectively, with a total of galaxies.
@@
The SN feedback model in SPHINX is calibrated to match the UV luminosity function (UVLF) at (see App. C in Rosdahl et al., 2018). A possible caveat is that the stellar mass-halo mass relation in SPHINX is – dex higher than predicted by, e.g., the Behroozi et al. (2019) model that matches a broad array of observational data including stellar mass functions at – and median UV—stellar mass relations at – (see Fig. 4 in Katz et al., 2023). Furthermore, the stellar mass function (SMF) in SPHINX is significantly overestimated compared to the SMF inferred from the JWST/NIRCam observations at – (Weibel et al., 2024). Hence, if stellar masses of SPHINX galaxies are overestimated, the amount of dust in the simulation should be higher than in the real Universe in order to match the UVLF. The overproduction of dust would then explain the fact that SPHINX predicts lower and luminosities at fixed stellar mass than those measured by JWST (e.g., Covelo-Paz et al., 2025; Meyer et al., 2024). Although the possibility of SPHINX galaxies having too much dust does not affect our main results as they are based on the intrinsic properties (Sect. 3), it will become important later for understanding the errors in the SFR() calibrations that use the observed (Sect. 4).
@@ 3 Results
In this section, we investigate the intrinsic (i.e., without taking into account dust attenuation) SFR— relation in the SPHINX20 cosmological simulation. We compare this relation to the SFR() calibrations from the literature and study how the variations in the physical properties of individual galaxies (metallicity, stellar age, etc.) impact the accuracy of the predicted SFRs. Based on this analysis, we develop a set of new SFR() calibrations that minimize the error in -derived SFRs at high redshifts.
Unless otherwise specified, we use the SFRs averaged over the last Myr of the star-formation history. This choice is motivated by the fact that nebular emission lines such as trace stars with lifetimes of –10 Myr (e.g., Kennicutt & Evans, 2012). This is further supported by the simulation itself, where the correlation coefficient between the SFR and peaks between Myr (–; see App. A).
@@ 3.1 The SFR— relation in SPHINX
Figure 1 shows the intrinsic SFR— relation for the galaxy sample selected from the SPHINX20 simulation as described in Sect. 2. In the same figure, we show different SFR() calibrations proposed in the literature. These include the original K98 calibration (), the K98 calibration converted to the Kroupa IMF (Kroupa, 2001b; ), and two calibrations from Theios et al. (2019, hereafter T19) calculated for a Kroupa-type IMF and solar metallicity (), and the same IMF but subsolar metallicity (; ), respectively. Here, is the conversion factor between and SFR, defined as follows:
@@
Figure 1 reveals that a constant cannot provide an accurate estimate of the SFRs of SPHINX galaxies across the full range. Specifically, the T19 calibration that assumes (the dashed blue line in Fig. 1) matches well the median SFR— relation in SPHINX at high luminosities (). In contrast, the same SFR() calibration at low luminosities () overestimates the SFRs by – dex. Decreasing would result in a better match at low luminosities; however, the SFRs at high luminosities would be underestimated in that case.
@@
We note that the median SFR— relation in SPHINX at low luminosities lies between the two T19 calibrations that differ only in the assumption on metallicity ( vs. ; indicated by the solid and dashed blue lines in Fig. 1, respectively). Motivated by this result, we test whether the variations in metallicity could be a key factor driving the evolution of the median SFR— relation in SPHINX. We split our sample into low- and high- groups ( and , respectively), and calculate the median stellar metallicity222Hereafter, stellar metallicities (and ages) of SPHINX galaxies are weighted by the LyC luminosity. in each group. We obtain and in the low- and high- groups, respectively. In other words, brighter galaxies are, on average, more metal-rich, which reflects the mass-metallicity relation (MZR) in SPHINX (for discussion, see Katz et al., 2023), and likely causes an upturn of the median SFR— relation at high luminosities (see also Fig. 1 where markers are color-coded by metallicity).
@@
Figure 1 further demonstrates that the original K98 calibration (indicated by the dotted orange line) overestimates the SFRs by – dex compared to SPHINX. One of the main reasons for this discrepancy is the difference in the IMFs: K98 uses the Salpeter (1955) IMF, whereas SPHINX uses a Kroupa-type IMF. After rescaling to the Kroupa IMF, the K98 calibration still predicts – dex higher SFRs than the simulation (the dashed orange line in Fig. 1). This difference cannot be eliminated even when using the T19 calibration which, unlike the K98 calibration, assumes the same stellar population synthesis model as SPHINX (i.e., BPASSv; the dashed blue line in Fig. 1). Therefore, the only way to fully reconcile the K98 calibration and the simulation data is to assume a sub-solar metallicity, for some galaxies as low as (the solid blue line in Fig. 1). As discussed in the introduction, such sub-solar metallicity is well in line with the current estimates of the gas-phase metallicity, as well as with the direct estimates of stellar metallicities of galaxies at (Steidel et al., 2016; Cullen et al., 2019), although we caution that those have been derived assuming constant star formation histories which might impact the results (Matthee et al., 2022).
@@
Figure 1 also reveals that the SFR— relation in SPHINX exhibits significant scatter around the median relation (hereafter ). In particular, we find that decreases from dex at to dex at , and has a median dex. This scatter is neglected by the existing SFR() calibrations, leading to errors in the predicted SFRs of individual galaxies. To quantify this effect, we introduce the SFR bias factor, defined as the ratio between the SFR predicted from and the “true” SFR measured in the simulation: . Here, we use the T19 calibration that assumes to calculate SFR() as this calibration provides the best match to the SPHINX data (see Fig. 1). We note that using a different calibration would only change by a constant factor (in logarithmic scale) without affecting any trends observed between and the physical properties of SPHINX galaxies.
@@
We find that decreases with stellar metallicity (; Fig. 2, top panel), in agreement with our earlier analysis of the median SFR— relation in SPHINX. However, we also report a significant scatter in at fixed : dex. This indicates that metallicity is not the only parameter responsible for the variations in the SFRs of SPHINX galaxies with respect to the SFRs predicted from . In particular, we find that shows a negative correlation with stellar age (; Fig. 2, bottom panel), with an average scatter of dex. High (– dex) that is found in galaxies with young stellar populations ( Myr) can be explained by the fact that the existing SFR() calibrations assume constant star formation over a long timescale (typically Myr or Gyr; e.g., Hao et al., 2011). This assumption breaks down in the simulation, where galaxies exhibit very bursty and stochastic star formation histories (Katz et al., 2023).
@@
We also examine the relationship between and stellar mass and find a negative correlation (), likely driven by the positive scaling of stellar mass with metallicity (see Katz et al., 2023 for the discussion of the MZR in SPHINX). Finally, we check the dependence of on the escape fraction of ionizing radiation () and find a nearly flat trend (), indicating no significant correlation. This result is expected since the strength of nebular emission lines is inversely proportional to the escape fraction of ionizing photons, , and SPHINX galaxies have, on average, low (%). This is in agreement with high-redshift observations, where is also estimated to be low ( %; e.g. Mascia et al., 2024).

### v2

# as a tracer of star formation in the SPHINX cosmological simulations

@@
as a tracer of star formation in the SPHINX cosmological simulations
The emission line in galaxies is a powerful tracer of their recent star formation activity. With the advent of JWST, we are now able to routinely observe in galaxies at high redshift () and thus measure their star formation rates (SFRs). However, using classical SFR() calibrations to derive the SFRs leads to biased results because high-redshift galaxies are commonly characterized by low metallicities and bursty star formation histories, affecting the conversion factor between the luminosity () and the SFR. In this work, we develop a set of new SFR() calibrations that allow us to predict the SFRs of -emitters at with minimal error. We use the SPHINX cosmological simulations to select a sample of star-forming galaxies representative of the -emitter population observed with JWST. We then derive linear corrections to the classical SFR() calibrations, taking into account variations in the physical properties (e.g., stellar metallicities) among individual galaxies. We obtain two new SFR() calibrations that, compared to the classical calibrations, reduce the root mean squared error (RMSE) in the predicted SFRs by dex and dex, respectively. Using the recent JWST NIRCam/grism observations of -emitters at , we show that the new calibrations affect the high-redshift galaxy population statistics: (i) the estimated cosmic star formation rate density decreases by %, and (ii) the observed slope of the star formation main sequence increases by .
@@
One of the most fundamental and basic properties of galaxies is the rate at which new stars form, the star formation rate (SFR). The SFR is sensitive to the interplay between the gas inflow and outflow rates of galaxies, and the efficiency at which this gas is converted into stars (e.g., Bouché et al., 2010; Lilly et al., 2013; Behroozi et al., 2019). The SFR can fluctuate on different timescales (e.g., Sparre et al., 2015; Iyer et al., 2020; Flores Velázquez et al., 2021), which range from dynamical timescales of molecular clouds (s of Myrs), through galaxy merger timescales ( Myr), to long (Gyr-scale) coherences in accretion rates (e.g., Matthee and Schaye, 2019; Wang and Lilly, 2020; Tacchella et al., 2020; Wan et al., 2025). Observationally, SFRs can be measured using various diagnostics (e.g., Kennicutt and Evans, 2012; De Looze et al., 2014; Davies et al., 2019) that are either based on the light from massive, short-lived stars (e.g., the rest-frame UV), or processed radiation (e.g., nebular lines in HII regions or dust emission). Each of these diagnostics probes different timescales of star formation and is affected by dust attenuation differently.
@@
At high redshifts (), the rest-frame UV emission used to be the most sensitive SFR indicator that could be measured for large galaxy samples, tracing star formation on Myr timescales (e.g., Bouwens et al., 2015). However, dust corrections in the rest-frame UV are significant, with potentially up to 50% of the cosmic star formation rate density () at being obscured by dust (e.g., Zavala et al., 2021; Sun et al., 2025). This emphasizes the great importance of alternative SFR indicators that are less sensitive to dust attenuation, such as the emission. The photons are emitted in the rest-frame optical ( Å), where dust corrections are smaller than in the rest-frame UV due to the negative slope of the dust attenuation curve (however, we caution that there is growing evidence of flatter attenuation curves at fixed or stellar mass at high redshifts compared to the local Universe; e.g., Shivaei et al., 2025; Fisher et al., 2025). The luminosity of the line is directly proportional to the luminosity of ionizing photons that are predominantly emitted by very young stellar populations, making sensitive to star formation on shorter timescales ( Myr; e.g., Leitherer et al., 1999). As the JWST can now measure the distribution of luminosities out to based on spectroscopy and narrow-band imaging (e.g., Clarke et al., 2024; Pirie et al., 2025; Covelo-Paz et al., 2025; Fu et al., 2025), it is timely to reassess the role of the H emission as a tracer of the SFR.
@@
The conversion factor between the H luminosity111Throughout this work, we assume a negligible contribution of active galactic nuclei to the H emission. and the SFR primarily depends on assumptions on the properties of stellar populations, in particular, the star formation history (SFH), the initial mass function (IMF), binarity, and stellar metallicity (e.g, Wilkins et al., 2019). The SFH and IMF are relevant because the ionizing photon emissivity from a single stellar population drops sharply after a few Myr when massive, short-lived O stars leave the main sequence. Binary interactions can extend these ionizing lifetimes (Götberg et al., 2017), smoothing the otherwise steep drop-off after a few Myrs. At fixed stellar mass and age, the metallicity affects the effective temperature of stellar atmospheres due to line blanketing effects, affecting the ionizing luminosity as well.
@@
The classical SFR() calibration proposed by Kennicutt (1998) (see also Kennicutt and Evans 2012) assumes a constant SFH over Myr and solar metallicity, based on the typical conditions in galaxies in the low-redshift Universe. However, various spectroscopic observations of galaxies at high redshifts suggest that their SFHs are bursty (e.g., Asada et al., 2024; Endsley et al., 2024; Cole et al., 2025; Carvajal-Bohorquez et al., 2025). Further, given the gas-phase oxygen abundances are typically measured to be % solar (Curti et al., 2024; Scholte et al., 2025), and early galaxies are expected to be -enhanced due to the rapid formation timescales (e.g., Chruślińska et al., 2024), the iron-peak element abundances are likely to be low. This implies hotter stars and a more efficient photon production at a fixed SFR (e.g., Theios et al., 2019) than at low redshift.
@@
The aim of this paper is to revisit the classical SFR() calibrations to predict the SFRs of galaxies at high redshift (). Instead of using idealistic models with constant SFHs and fixed stellar metallicities, we use simulated galaxies and their modeled luminosities from the SPHINX cosmological radiation-hydrodynamical simulation (Rosdahl et al., 2018; Katz et al., 2023). These cosmological simulations self-consistently model SFHs and metal enrichment processes and are calibrated to match various observables at high redshifts (). The simulated volume matches the typical volumes of deep JWST surveys well, which enables a high resolution that can capture SFR variations on short timescales ( Myr) and approximate the distribution of gas in the interstellar medium relevant for radiative transfer effects.
@@
In Sect. 2, we briefly describe the SPHINX simulations and explain how the H luminosities are derived. In Sect. 3, we present a set of simulation-inferred SFR() calibrations that predict the SFRs of SPHINX galaxies with minimal error. Finally, we discuss the implications of our work in Sect. 4 and summarize in Sect. 5.
@@ 2 Methods
In this work, we use the simulated galaxy data from SPHINX, a suite of cosmological radiation-hydrodynamic simulations (Rosdahl et al., 2018). SPHINX leverages the adaptive mesh refinement code RAMSES-RT (Rosdahl et al., 2013; Teyssier, 2002) to simulate tens of thousands of galaxies in the epoch of reionization. The ISM is resolved down to co-moving pc, corresponding to a physical resolution of pc at . The dark matter halos are resolved down to the atomic cooling threshold ( M⊙). Hydrodynamics, non-equilibrium thermochemistry, and radiative transfer of Lyman-continuum radiation (LyC; Å) are performed self-consistently on the fly, with subgrid implementations for gas cooling, star formation, metal enrichment, and feedback. The on-the-fly radiative emission and propagation of LyC photons make SPHINX an excellent model for accurately predicting how emission traces star formation in the simulated galaxies.
@@
Our analysis is based on the data from the SPHINX20 Public Data Release, Version 1 (SPDRv1; Katz et al., 2023). SPHINX20 is the largest simulation of the SPHINX suite, with a periodic box with volume of ( cMpc)3 simulated down to a redshift of (Rosdahl et al., 2022). SPHINX20 uses the BPASSv stellar population synthesis (SPS) model (Stanway and Eldridge, 2018), which adopts a Kroupa (2001) IMF with an upper mass cutoff of 100 M⊙ and a slope of from to M⊙ and from to M⊙. The intrinsic Balmer line luminosities are generated by interpolating the CLOUDY v17.03 models (Ferland et al., 2017) for gas cells that host stellar particles with unresolved Stromgren spheres, or by using the non-equilibrium ionizing fractions from the simulation otherwise. For each halo, the total intrinsic line emission is then the sum of the line luminosities of all gas cells within the virial radius of the halo. The equivalent widths are calculated by taking the stellar and nebular continuum emission into account, generated as detailed in Katz et al. (2023).
We post-process the line and continuum radiation, subject to absorption and scattering by dust, using the radiative transfer code RASCAS (Michel-Dansac et al., 2020). Dust is assigned to the gas cells using the Small Magellanic Cloud (SMC) dust model from Laursen et al. (2009). We note that the dust attenuation curve at high redshifts likely has a steeper slope than the SMC slope (e.g., Shivaei et al., 2025; Markov et al., 2025), possibly due to changes in the dust grain properties. Next, 200,000 photon packets are probabilistically distributed to the gas cells, with the initial wavelengths placed at the line centers, thermally broadened, and shifted according to the bulk velocity of the cell. Likewise, 200,000 photon packets are probabilistically distributed to the star particles (gas cells) based on the luminosities of the stellar (nebular) continuum at the line centers. The observed line luminosities and equivalent widths are then obtained by computing the escape fractions of each component (for details, see Katz et al., 2023).
@@
Due to the computation time and data storage limitations, SPDRv1 only includes SPHINX20 galaxies with M⊙ yr-1, where is the SFR averaged over Myr. This excludes objects that are likely too faint to be observed with JWST (Katz et al., 2023), unless lensing is considered (e.g., Bezanson et al., 2024; Naidu et al., 2024). However, an SFR cut also leads to an incompleteness in the SFR— relation at low luminosities, which can bias our results. We therefore limit our selection of galaxies from SPDRv1 to those with an observed luminosity () greater than erg s-1 (here, we use the simulation-based luminosities calculated as described above). This value is motivated by the recent LF measurements from JWST NIRCam/grism, which reach similar luminosities at the faint end (e.g., Fu et al., 2025). Crucially, this selection yields a complete -emitter sample, which we check by running RASCAS on all SPHINX20 halos whose intrinsic luminosity () exceeds the luminosity cut. We also try erg s-1 and erg s-1 cuts, finding little impact on our results. With this additional luminosity cut, our final sample comprises , 66, 48, 36, 22, 17, and 13 galaxies at , 5, 6, 7, 8, 9, and 10, respectively, with a total of galaxies.
@@
The SN feedback model in SPHINX is calibrated to match the UV luminosity function (UVLF) at (see App. C in Rosdahl et al., 2018). A possible caveat is that the stellar mass-halo mass relation in SPHINX is – dex higher than predicted by the Behroozi et al. (2019) model, for example, which matches a broad array of observational data, including stellar mass functions at – and median UV—stellar mass relations at – (see Fig. 4 in Katz et al., 2023). Furthermore, the stellar mass function (SMF) in SPHINX is significantly overestimated compared to the SMF inferred from the JWST/NIRCam observations at – (Weibel et al., 2024). If the stellar masses of SPHINX galaxies are overestimated, the SFRs should also be overestimated on average. This implies that the amount of dust in the simulation should be higher than in the real Universe in order to match the UVLF. The excess of dust would then explain the fact that SPHINX predicts lower and luminosities at fixed stellar mass than those measured by JWST (e.g., Covelo-Paz et al., 2025; Meyer et al., 2024). Although the possibility of SPHINX galaxies having too much dust does not affect our main results as they are based on the intrinsic properties (Sect. 3), it will become important later for understanding the errors in the SFR() calibrations that use the observed (Sect. 4).
@@ 3 Results
In this section, we investigate the intrinsic (i.e., without taking dust attenuation into account) SFR— relation in the SPHINX20 cosmological simulation. We compare this relation to the SFR() calibrations from the literature and study the effect of the variations in the physical properties of individual galaxies (metallicity, stellar age, etc.) on the accuracy of the predicted SFRs. Based on this analysis, we develop a set of new SFR() calibrations that minimize the error in -derived SFRs at high redshift.
Unless otherwise specified, we use the SFRs averaged over the last Myr of the SFH. This choice is motivated by the fact that nebular emission lines such as trace stars with lifetimes of –10 Myr (e.g., Kennicutt and Evans, 2012). This is further supported by the simulation itself, where the correlation coefficient between the SFR and peaks between Myr (–; see App. A).
@@ 3.1 The SFR— relation in SPHINX
Figure 1 shows the intrinsic SFR— relation for the galaxy sample selected from the SPHINX20 simulation as described in Sect. 2. In the same figure, we show different SFR() calibrations proposed in the literature. These include the original K98 calibration (), the K98 calibration converted to the Kroupa IMF (Kroupa, 2001; ), and two calibrations from Theios et al. (2019, hereafter T19) calculated for a Kroupa-type IMF and solar metallicity (), and the same IMF but subsolar metallicity (; ), respectively. Here, is the conversion factor between and SFR, defined as follows:
@@
Figure 1 reveals that a constant cannot provide an accurate estimate of the SFRs of SPHINX galaxies across the full range. Specifically, the T19 calibration that assumes (the dashed blue line in Fig. 1) matches the median SFR— relation in SPHINX at high luminosities well (). In contrast, the same SFR() calibration at low luminosities () overestimates the SFRs by – dex. Decreasing would result in a better match at low luminosities, but the SFRs at high luminosities would then be underestimated.
@@
We note that the median SFR— relation in SPHINX at low luminosities lies between the two T19 calibrations that differ only in the assumption on metallicity ( vs. ; indicated by the solid and dashed blue lines in Fig. 1, respectively). Motivated by this result, we test whether the variations in metallicity might be a key factor driving the evolution of the median SFR— relation in SPHINX. We split our sample into low- and high- groups ( and , respectively), and calculate the median stellar metallicity222Hereafter, stellar metallicities (and ages) of SPHINX galaxies are weighted by the LyC luminosity. in each group. We obtain and in the low- and high- groups, respectively. In other words, brighter galaxies are more metal-rich on average, which reflects the mass-metallicity relation (MZR) in SPHINX (for a discussion, see Katz et al., 2023), and likely causes an upturn of the median SFR— relation at high luminosities (see also Fig. 1, where the markers are color-coded by metallicity).
@@
Figure 1 further demonstrates that the original K98 calibration (indicated by the dotted orange line) overestimates the SFRs by – dex compared to SPHINX. One of the main reasons for this discrepancy is the difference in the IMFs: K98 uses the Salpeter (1955) IMF, whereas SPHINX uses a Kroupa-type IMF. After rescaling to the Kroupa IMF, the K98 calibration still predicts – dex higher SFRs than the simulation (the dashed orange line in Fig. 1). This difference cannot be eliminated even when using the T19 calibration which, unlike the K98 calibration, assumes the same stellar population synthesis model as SPHINX (i.e., BPASSv; the dashed blue line in Fig. 1). Therefore, the only way to fully reconcile the K98 calibration and the simulation data is to assume a sub-solar metallicity, for some galaxies as low as (the solid blue line in Fig. 1). As discussed in the introduction, a subsolar metallicity like this agrees well with the current estimates of the gas-phase metallicity and also with the direct estimates of stellar metallicities of galaxies at (Steidel et al., 2016; Cullen et al., 2019), although we caution that these were derived assuming constant SFHs, which might impact the results (Matthee et al., 2022).
@@
Figure 1 also reveals that the SFR— relation in SPHINX exhibits significant scatter around the median relation (hereafter, ). In particular, we find that decreases from dex at to dex at , and has a median dex. This scatter is neglected by the existing SFR() calibrations, leading to errors in the predicted SFRs of individual galaxies. To quantify this effect, we introduce the SFR bias factor, defined as the ratio of the SFR predicted from to the true SFR directly tracked by the simulation: . Here, we use the T19 calibration that assumes to calculate SFR() as this calibration provides the best match to the SPHINX data (see Fig. 1). We note that using a different calibration would only change by a constant factor (in logarithmic scale) without affecting any trends observed between and the physical properties of SPHINX galaxies.
@@
We find that decreases with stellar metallicity (, ; Fig. 2, top panel), in agreement with our earlier analysis of the median SFR— relation in SPHINX. However, we also report a significant scatter in at fixed : dex. This indicates that metallicity alone does not cause the variations in the SFRs of SPHINX galaxies with respect to the SFRs predicted from . In particular, we find that shows a negative correlation with stellar age (, ; Fig. 2, bottom panel), with an average scatter of dex. Notably, galaxies with young stellar populations ( Myr) exhibit high (– dex). Even if we use the best-fit SFR() calibration in the form of Eq. 1 (), still reaches dex at – Myr. This bias can be explained by the fact that the existing SFR() calibrations assume constant star formation over a long timescale (typically Myr or Gyr; e.g., Hao et al., 2011). This assumption breaks down in the simulation, where galaxies exhibit very bursty and stochastic SFHs (Katz et al., 2023).
@@
We also examine the relationship between and stellar mass and find a negative correlation (, ), likely driven by the positive scaling of stellar mass with metallicity (see Katz et al., 2023 for the discussion of the MZR in SPHINX). Finally, we check the dependence of on the escape fraction of ionizing radiation () and find a nearly flat trend (, ). This result is expected because the strength of nebular emission lines is inversely proportional to the escape fraction of ionizing photons, , and SPHINX galaxies have low (%) on average. This agrees with high-redshift observations, where is also estimated to be low (%; e.g., Mascia et al., 2024).